# Masked Mixers for Language Generation and Retrieval

### v2

@@ 6.3 Masked Mixers are more suitable for cosine similarity InfoNCE -trained retrieval
The two major downsides to the secondary retrieval model training approach is that there is no guarantee that the embedding model is capable of capturing the information necessary to perform retrieval, as the embedding model was trained for an entirely different task (causal language modeling in this case), and that inference requires much more computation. Empirically the first downside is observed to be the case as the retrieval model tends to overfit the retrieval dataset. It is therefore pertinent to observe the ability of the transformer and masked mixer to be trained to form embedding suitable for retrieval and one particularly promising way to do this is to perform noise contrastive training on a causal language-pretrained model.
@@
We implement a temperature-scaled cosine distance-based InfoNCE loss where $q^{+}$ indicates the query embedding, $\tau=0.02$ the temperature hyperparameter set to 0.02, $d^{+}$ the matching text embedding and $d_{i}$ randomly sampled nonmatching text embeddings (Equation (7)).
@@
We use the model’s last hidden layer of the second-to-last token as the embedding as shown in Equation 8 instead of appending an extra token to each sequence, and sample with a context window size of 512. We include no hard negatives and sample one query with one matching and thirty non-matching text segments, and process four batches per optimizer update (one per GPU). We sample each batch using an analogous method to Algorithm 1 except that token sequences rather than embeddings are retrieved, and each input is assembled in the batch dimension such that a batch size of 32 has 32 separate text sequences.
@@
We first benchmarked a state-of-the-art (when this work was commenced) large model pretrained on billions of tokens and postrained via cosine similarity InfoNCE for retrieval (among other tasks), e5 Mistral Instruct [32]. When we applied to the synthetic Fineweb retrieval dataset introduced in the last section we observed that top-1 evaluation accuracy for 32- and 64-sample retrieval of around 100% for this model (compared to 92.8% and 87.7% respectively for the masked mixer as trained using the method above) suggesting that this dataset is not particularly challenging for a strong model like e5 Mistral Instruct. We therefore generated a new synthetic dataset in which the same Llama model was instructed to generate shorter queries for FineMath 4+ text segments. This can be viewed as being a more difficult retrieval task because the dataset itself is restricted in subject and also because the queries are themselves less shorter and descriptive than they were for the FineWeb.
The Top-1 retrieval accuracy of each model with the FineMath 4+ evaluation Cross-Entropy Loss for models trained on this dataset are shown in Table 2. For both top-1 retrieval accuracy and loss values, measurements were performed on evaluation datasets held out during each respective training, with the pretraining corpus including both training and evaluation target text segments used for retrieval and none of the queries. All models except the already-trained e5 Mistral Instruct were trained using n=180k retrieval samples (one sample being a query with its matching target text) except where noted, and evaluated on a hold-out dataset of 20k samples (n=400k implies 380k training and 20k evaluation samples). Note that e5 Mistral Instruct uses a different tokenizer such that its cross-entropy loss is not comparable, and that due to the masked mixer’s far fewer inter-token parameters the transformer in the first line of Table 2 was pretrained using roughly the same compute as the $n_{l}=32$ masked mixer and $d_{m}=1024$ masked mixer and around double the compute as the masked mixer in the second line of that table.
@@
Table 2: FineMath evaluation retrieval top-1 accuracy with n=32 sample size per retrieval, postrained on s=180k synthetic retrieval samples with $d_{m}=512$ except where denoted.
For practical query retrieval tasks one usually wants to perform retrieval using sample sizes of hundreds or thousands of targets. This is a major drawback for benchmark datasets which compare queries to only a few targets (for example as is typical for datasets like MSMarco [33]) because few-sample retrieval performance has no guarantee of extrapolating to many-sample retrieval. It is therefore important to measure many-sample retrieval performance directly, and the these results are shown in Tables 1 and 3. Most notably, the transformer model’s near-parity performance to the much larger e5 Mistral Instruct on small-context inputs vanishes as the sample size increases such that it becomes far worse for large-context retrieval relative to e5 Mistral, whereas the masked mixer (trained on identical data) exhibits uniformly superior performance to the other models across all sample sizes. As has been observed for transformers, increasing the masked mixer’s pretraining compute increases the resulting model’s retrieval performance (Table 3), and pretraining on an I.I.D. dataset to the retrieval task yields increased re trieval performance than pretraining on a dataset with a mix of in- and out of- distribution data such as that in the FineWeb (Table S7).
@@ 6.4 Autoencoders yield effective pretrained embeddings for InfoNCE
We conclude that autoencoding provides a promising new direction for pretraining for retrieval models as it appears to be far less sensitive to the amount of compute used, although more work needs to be done to more fully determine the strengths and weaknesses of this approach as compared to causal language modeling-based pretraining.
@@ 7 Limitations of this work
This study is limited in scale: experiments were performed using a relatively small datasets (none larger than 10 billion tokens) and compute (four V100 GPUs) and model sizes (tens or hundreds of millions of parameters). Although there are reasons to suspect that the causal language modeling results here would translate to much larger datasets and compute, this has yet to be investigated as training runs were performed in 30h or less on 4x V100s (for a maximum theoretical limit of $5.4*10^{19}$ Tensor FLOPs), which is a relatively small amount by modern standards. Another limitation inherent in this work is that compute efficiency is measured in a concrete hardware-specific fashion. We chose this route as it is immediately applicable to the primary goal of language model training (ie getting the lowest possible loss in the least amount of time), and one can recover the FLOPS per experiment using available data on each compute node.
A downside to our wall clock and hardware-specific performance measurements is that our GPUs are not typically used for language model training in industrial settings today, where H100 or A100s are favored. Newer GPUs are specially tailored for transformer inference and training and thus our comparisons between transformer and mixer architectures would likely be somewhat different if performed on these newer GPUs. H100s in particular have dedicated transformer engines [34], which would be expected to provide speedups for that model relative to other architectures.
Our input representation methods yield results that are consistent with other invertibility investigations (autoencoding performance for example) but are by no means comprehensive. It is entirely possible that a different input representation method would give more accuracy to trained transformers or less to masked mixers.
@@ 8 Discussion
8.1 Input representation Accuracy and Model Efficiency: What Good is Invertiblity for Deep Learning?
This work has explored the use of input representation accuracy as a guide to improving deep learning model architectures for language generation and retrieval, and has found that global invertibility is indeed very useful in estimating performance capabilities. This is not at all a surprising result when viewed from the theory of numerical analysis, as invertibility is usually the second-most important facet of a system (after linearity). It has also become is evident that there is no simple relationship between representation accuracy and either generation or retrieval or else very small masked mixers would outperform transformers, and untrained mixers would for effective embeddings for retrieval. It is evident that other considerations (for example, the set of possible transformations a model is capable) are important in varying degrees as well.
@@ 8.2 Semi-recurrent and feedforward architectures
Transformers blend features of recurrent neural networks and pure feedforward models: during training they require only one forward and backward pass per input for training on all tokens of that input (like feedforward models) but during inference they may be used for language generation of an indeterminate number of tokens (assuming that the positional encoding method used is sufficiently flexible). The main downside is that these models are effectively $O(n^{2}d)$ in both time and space complexity during training because for each next token for an input of $n$ tokens, the key and value projections must be re-computed. Masked mixers as implemented are pure feedforward models in the sense that these models have a fixed output size for both training and inference.
We can also comment on similarities and differences between the masked mixer and other language models replacing attention with specialized convolutions such as the Monarch mixer [14] or Hyena [15] architectures, which are motivated primarily by an effort to reduce the $O(n^{2}d)$ computational complexity inherent in the transformer architecture. Firstly, we note that although all three works are focused on increasing the efficiency of language modeling, our work has an entirely different motivation (increased representational efficiency rather than lower bounds for computational scaling) and thus focuses on training efficiency at the expense of inference efficiency. In particular, masked mixers have higher inference complexity ($O(n^{3}d)$) than transformers, which can cache hidden layer values. Caching masked mixer hidden state values during inference would be useful for convolutions with restricted substructure (for example, parameterized as Toeplitz or Monarch matrices) but we leave the investigation of such models to future work.
A major difference between this work and many comparisons of sequence-length efficient architectures to transformers in the literature is that the latter do not typically use transformers with the latest architectural features combined with extensive hyperparameter and model size optimizations when establishing a performance baseline for a GPT-style causal language model. There is often good reason for this: transformers have benefited from years of optimizations that often drastically improve training efficiency and new architectures have not had this kind of effort applied.
@@ 8.3 How much more efficient are masked mixers compared to transformers for retrieval?
The results in Section 6.3 indicate that masked mixers are substantially more effective retrieval models compared to transformers receiving the same training protocol, and we include results from the near-SOTA e5 Mistral instruct as a benchmark of what a very capable off-the-shelf model can do for this dataset. Finding exactly how much more efficient the masked mixer architecture is for retrieval would require knowledge of the scaling properties of each model type, work that we leave to future investigations.
@@
With this in mind we do not mean to imply that masked mixers are a factor of 10,000 more efficient than transformers for retrieval, only that a result achieved using a large transformer model may be surpassed with a small masked mixer using 1/10,000th the compute. We present evidence for the hypothesis that this results primarily from the use of masked mixers for retrieval and also from the training of this model on I.I.D. data to the evaluation set.
@@ 8.4 Is attention useful for efficient language modeling?
From the results presented in this paper it would appear that properly optimized attention mechanisms do confer benefits to the process of efficient causal language model training for larger context windows, at least until more efficient training can be obtained using attentionless models such as masked mixers. On the other hand the evidence suggests that attention is generally not desirable for language retrieval tasks, as is predicted by the model invertibility framework presented in this paper.
@@ 8.5 How many parameters are necessary for language modeling?
For more concrete estimates based on datasets used today, consider that language models today are nearly universally trained to predict every ‘next’ token in a concatenated sequence of tokens. This means that for a training dataset of $n$ tokens, the task is approximate $n$ points in some presumably very high-dimensional space. The FineWeb-edu 10BT dataset would thus require a model of dimension $1848\ln(10\times 10^{9})\approx 8*23=$ 184 for perfect memorization, and this number grows extremely slowly as the dataset size increases. The largest dataset open-source models are currently trained on is 15 trillion tokens (for Llama-3 [5]), and in this case we can represent this dataset arbitrarily well in a space of approximately $2408\ln(15\times 10^{12})\approx 8*30=$ 240 dimensional. The goal of language models is to generalize to larger datasets and thus we would hope the model would accurately predict the next tokens of a much larger dataset. But even supposing that this 15 trillion token dataset is only one millionth of the size of this generalized dataset, one would still only require a space of dimension $3528\ln(15\times 10^{18})\approx 8*44=$ 352.
How does this relate to the number of parameters necessary in a model? Two options are either it is equivalent to the dimensionality of the model with respect to the points it can approximate, or else the model’s ‘width’ or hidden layer dimension is equivalent to the model’s dimensionality. Because the vector space of successive layers of practically every modern deep learning model are dependent (activations in layer $n+1$ depend on layer $n$ and perhaps layer $n-1$ etc.) and dimensionality is defined on linear independence, it seems more likely that a model’s dimensionality best corresponds with its hidden layer dimension. If this is true then a model of no more than around 1300 hidden layer elements should be capable of completing any English language sentence, or a model with a width of 350 can accurately predict a next token for a massive dataset. If these hidden widths were used for default Llama architectures, the resulting models would are around 100 million (given a hidden width of $d_{m}=350$) and 995 million (for $d_{m}=1300$) parameters.

### v3

@@ 6.3 Masked mixers are more suitable for cosine similarity InfoNCE -trained retrieval
The two major downsides to the indirect retrieval model training approach is that there is no guarantee that the embedding model is capable of capturing the information necessary to perform retrieval as the embedding model was trained for an entirely different task (causal language modeling in this case), and that inference requires much more computation assuming that the retrieval model is composed of more than one matrix multiplication. Empirically the first is observed to be the case for FineWeb as the retrieval model tends to overfit the retrieval dataset. It is therefore pertinent to observe the ability of the transformer and masked mixer to be trained to form embedding suitable for retrieval and one particularly promising way to do this is to perform noise contrastive training on a causal language-pretrained model.
We implement a temperature-scaled cosine distance-based InfoNCE loss where $q^{+}$ indicates the query embedding, $\tau=0.02$ the temperature hyperparameter set to 0.02, $d^{+}$ the matching text embedding and $d_{i}$ randomly sampled non-matching text embeddings as shown in Equation (7), after [33].
@@
For (7) we use the model’s last hidden layer of the second-to-last token as the embedding for cosine similarity computation shown in Equation (8) instead of appending an extra token to each sequence, and sample with a context window size of 512. We include no hard negatives and sample one query with one matching and thirty non-matching text segments, and process four batches per optimizer update (one per GPU). We sample each batch using an analogous method to Algorithm 1 except that token sequences rather than embeddings are retrieved, and each input is assembled in the batch dimension such that a batch size of 32 has 32 separate text sequences.
@@
We first benchmarked a state-of-the-art (when this work was commenced) large model pretrained on trillions of tokens and postrained via cosine similarity InfoNCE for retrieval (among other tasks), e5 Mistral Instruct [33]. When we applied to the synthetic Fineweb retrieval dataset introduced in the last section we observed that top-1 evaluation accuracy for 32- and 64-sample retrieval of around 100% for this model (compared to 92.8% and 87.7% respectively for the masked mixer as trained using the method above) suggesting that this dataset is not particularly challenging for a strong model like e5 Mistral Instruct. We therefore generated a new synthetic dataset in which the same Llama model was instructed to generate shorter queries for FineMath 4+ text segments. This can be viewed as being a more difficult retrieval task because the dataset itself is restricted in subject and also because the queries are themselves less shorter and descriptive than they were for the FineWeb.
The top-1 retrieval accuracy of each model with the FineMath 4+ evaluation Cross-Entropy Loss for models trained on this dataset are shown in Table 2. For both top-1 retrieval accuracy and loss values, measurements were performed on evaluation datasets held out during each respective training, with the pretraining corpus including both training and evaluation target text segments used for retrieval and none of the queries. All models except the already-trained e5 Mistral Instruct were trained using n=180k retrieval samples (one sample being a query with its matching target text) except where noted, and evaluated on a hold-out dataset of 20k samples. Note that e5 Mistral Instruct uses a different tokenizer such that its cross-entropy loss is not comparable, and that due to the masked mixer’s far fewer inter-token parameters the transformer in the first line of Table 2 was pretrained using roughly the same compute as the $n_{l}=32$ and $d_{m}=1024$ masked mixers and around double the compute as the masked mixer in the second line of that table.
@@
Table 2: FineMath evaluation retrieval top-1 accuracy with n=32 sample size per retrieval, postrained on s=180k synthetic retrieval samples with $n_{l}=16$ and $d_{m}=512$ except where denoted.
For practical query retrieval tasks such as large scale RAG, one usually wants to perform retrieval using sample sizes of hundreds or thousands of targets. This is a major drawback for benchmark datasets which compare queries to only a few targets (for example as is typical for datasets like MSMarco [34]) because few-sample retrieval performance has no guarantee of extrapolating to many-sample retrieval. It is therefore important to measure many-sample retrieval performance directly, and the these results are shown in Tables 1 and 3. Most notably, the transformer model’s near-parity performance to the much larger e5 Mistral Instruct on small-context inputs vanishes as the sample size increases such that it becomes far worse for large-context retrieval relative to e5 Mistral, whereas the masked mixer (trained on identical data) exhibits uniformly superior performance to the other models across all sample sizes.
@@
As has been observed for transformers, increasing the masked mixer’s pretraining compute increases the resulting model’s retrieval performance (Table 3), and pretraining on an I.I.D. dataset to the retrieval task yields increased retrieval performance than pretraining on a dataset with a mix of in- and out of- distribution data such as that in the FineWeb (Table S7). These effects are dataset-dependent, however, as we find accurate FineWeb retrieval using models pre- and post-trained on FineMath 4+ (Table S8). For models trained on the FineWeb and evaluated on FineMath we find that the masked mixer is once again more accurate than the transformer, especially for large sample sizes (Table S8).
From the large increase in accuracy upon dataset size doubling, it may be wondered whether doubling the dataset again would allows the transformer to match the performance of the masked mixer with a smaller dataset. When this is done, we find that the transformer gains little accuracy and is not near matching the masked mixer’s accuracy, particularly for large-sample retrieval (Table 4). A preliminary scaling study of top-1 accuracies achieved for transformers and masked mixers postrained with various dataset sizes (for retrieval with $s=8192$) is shown in Figure S14, and demonstrates that one would require many orders of magnitude more retrieval training data for the transformer to match the mixer’s accuracy.
Model $n=32$ 64 128 256 512 1024 2048 4096 8192
Transformer 95.8 93.3 90.2 86.0 80.8 74.4 67.6 60.5 52.6
Masked Mixer 98.6 97.8 96.5 94.5 91.7 88.5 83.8 79.0 73.7
Table 4: FineMath 4+ evaluation retrieval top-1 accuracy (%) for various sample sizes $n$, post-trained on s=780k synthetic retrieval samples
@@ 6.4 Autoencoders yield effective pretrained embeddings for InfoNCE
We conclude that autoencoding provides a promising new direction for pretraining for retrieval models as this pretraining method appears to be far less sensitive to the amount of compute used, although more work needs to be done to more fully determine the strengths and weaknesses of this approach as compared to causal language modeling-based pretraining.
@@ 7 Limitations Of This Work
This study is limited in scale: experiments were performed using a relatively small datasets (none larger than 10 billion tokens) and compute (four V100 GPUs) and model sizes (tens or hundreds of millions of parameters). Although there are reasons to suspect that the causal language modeling results here would translate to much larger datasets and compute, this has yet to be investigated as training runs were performed in 30h or less on 4x V100s (for a maximum theoretical limit of $5*10^{19}$ Tensor FLOPs), which is a relatively small amount by modern standards. Another limitation inherent in this work is that compute efficiency is measured in a concrete hardware-specific fashion. We chose this route as it is immediately applicable to the primary goal of language model training (ie getting the lowest possible loss in the least amount of time), and one can recover the FLOPS per experiment using available data on each compute node.
A downside to our wall clock and hardware-specific performance measurements is that our GPUs are not typically used for language model training in industrial settings today, where H100 or A100s are favored. Newer GPUs are specially tailored for transformer inference and training and thus our comparisons between transformer and mixer architectures would likely be somewhat different if performed on these newer GPUs. H100s in particular have dedicated transformer engines [35], which would be expected to provide speedups relative to other architectures. Likewise, our approach used one set of software implementations (Pytorch, Huggingface Transformers, Accelerate etc.) and our conclusions with respect to training efficiencies are limited by the choices made in those libraries.
Our input representation methods yield results that are consistent with other invertibility investigations (autoencoding performance for example) but are by no means comprehensive. It is entirely possible that a different input representation method would yield different results, however.
@@ 8 Discussion
8.1 Input representation accuracy and training efficiency: what good is invertibility for deep learning?
This work has explored the use of input representation accuracy and global invertibility as a guide to improving deep learning model architectures for language generation and retrieval, and has found that global invertibility is indeed very useful in estimating performance capabilities. This is not at all a surprising result when viewed from the theory of numerical analysis, as invertibility is usually the second-most important facet of any numerical framework (after linearity). It has also become is evident that there is no simple relationship between representation accuracy and either generation or retrieval or else very small masked mixers would outperform transformers, and untrained mixers would for effective embeddings for retrieval. It is evident that other considerations (for example, the set of possible transformations a model is capable) are important in varying degrees as well.
@@ 8.2 Semi-recurrent and feedforward architectures
Transformers blend features of recurrent neural networks and pure feedforward models: during training they require only one forward and backward pass per input sequence to generate useful gradients using the loss from each output sequence element (like feedforward models) but during inference they may be used for language generation of an indeterminate number of tokens (assuming that the positional encoding method used is sufficiently flexible). The main downside is that these models are effectively $O(n^{2}d)$ in both time and space complexity during training because for each next token for an input of $n$ tokens, the key and value projections must be re-computed. Masked mixers as implemented are pure feedforward models in the sense that these models have a fixed output size for both training and inference.
We can also comment on similarities and differences between the masked mixer and other language models replacing attention with specialized convolutions such as the Monarch mixer [14] or Hyena [15] architectures, which are motivated primarily by an effort to reduce the $O(n^{2}d)$ computational complexity inherent in the transformer architecture. Our work has an entirely different motivation, namely increased representational efficiency rather than lower bounds for computational scaling, and thus focuses on training efficiency at the expense of inference efficiency. In particular, masked mixers have higher inference complexity ($O(n^{3}d)$) than transformers, which can cache hidden layer values. Caching masked mixer hidden state values during inference would be useful for convolutions with restricted substructure (for example, parameterized as Toeplitz or Monarch matrices) but we leave the investigation of such models to future work.
@@ 8.3 How much more efficient are masked mixers compared to transformers for retrieval?
The results in Section 6.3 indicate that masked mixers are substantially more effective retrieval models compared to transformers receiving the same training protocol, and we include results from the near-SOTA e5 Mistral instruct as a benchmark of what a very capable off-the-shelf model can do for this dataset.
@@
Another way to measure the relative efficiency of the transformer versus the mixer for retrieval is to find what amount of training data is necessary for one model’s performance to match the other’s. The transformer’s accuracy for $s=8192$ retrieval increases by 31.6% and 2.2% for two dataset doublings (200k to 400k to 800k total samples) suggest that one would need to transform this dataset size by $31.6/2.2>14$ logarithms for linear interpolation, yielding an impossibly large number of samples necessary for the transformer to match the masked mixer’s performance. If we relax this value by a large amount and instead interpolate once (such that every next dataset doubling yields $0.153$ percent accuracy increase, one requires a dataset $2^{(70.6-50.4)/0.153}\approx 2^{132}>5*10^{39}$ as large as that for the masked mixer.
Finding exactly how much more efficient the masked mixer architecture is for retrieval would require more knowledge of the scaling properties of each model type with respect to parameter number, pretraining dataset size, and post-training dataset size. We leave this work to future investigations.
@@ 8.4 Is attention useful for efficient language modeling?
In this work we show that as a general rule, attention is useful for language tasks that would not be expected to require a globally invertible function, but this transformation does not appear to be useful for tasks requiring invertibility (autoencoding and retrieval in particular).
@@ 8.5 How many parameters are necessary for language modeling?
For more concrete estimates, consider the causal language modeling datasets used to train models to predict every ‘next’ token in a concatenated sequence of tokens. This means that for a training dataset of $n$ tokens, the task is approximate $n$ points in some presumably very high-dimensional space. The FineWeb-edu 10BT dataset would thus require a model of dimension $1848\ln(10\times 10^{9})\approx 8*23=$ 184 for perfect memorization, and this number grows extremely slowly as the dataset size increases. One of the largest datasets that open-source models are currently trained on is 15 trillion tokens (for Llama-3 [5]), and in this case we can represent this dataset arbitrarily well in a space approximately $2408\ln(15\times 10^{12})\approx 8*30=$ 240 dimensional. The goal of language models is to generalize to larger datasets and thus we would hope the model would accurately predict the next tokens of a much larger dataset. But even supposing that this 15 trillion token dataset is only one millionth of the size of this generalized dataset, one would still only require a space of dimension $3528\ln(15\times 10^{18})\approx 8*44=$ 352.
How does this relate to the number of parameters necessary in a model? Parameter number may be equivalent to the dimensionality of the model with respect to the points it can approximate, or else the model’s ‘width’ or hidden layer dimension is equivalent to the model’s dimensionality. Because the vector space of successive layers of practically every modern deep learning model are dependent (activations in layer $n+1$ depend on layer $n$ and perhaps layer $n-1$ etc.) and dimensionality is defined on linear independence, it seems more likely that a model’s dimensionality best corresponds with its hidden layer dimension. If this is true then a model of no more than around 1300 hidden layer elements should be capable of completing any English language sentence, or a model with a width of 350 can accurately predict a next token for a massive dataset. If these hidden widths were used for default Llama architectures, the resulting models would are around 100 million (given a hidden width of $d_{m}=350$) and 995 million (for $d_{m}=1300$) parameters.